# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ A.6 Proofs Using Nonuniform Learnability
[[[[of Theorem A.14]]]] From Lemma A.16 we have that
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\tilde{\epsilon}_{T}\right)\geq 1-\delta\,.$ (23)
We can now use Lemma A.17 to obtain with probability of at least $1-\eta$ over the sampling of $h_{t}$ that $T\leq\frac{\log\left(2/\eta\right)}{\hat{p}}$ and therefore
$\tilde{\epsilon}_{T}$ $=\frac{\log T+2\log\log\left(T+1\right)+\log{4/\delta}}{N}$
$\leq\frac{\log\left(\frac{\log\left(2/\eta\right)}{\hat{p}}\right)+2\log\log\left(\frac{\log\left(2/\eta\right)}{\hat{p}}+1\right)+\log{4/\delta}}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}}\right)+\log\left(\frac{4}{\delta}\right)+\log\left(\log\left(\frac{2}{\eta}\right)\right)+2\log\log\left(\frac{\log\left(\frac{2}{\eta}\right)}{\hat{p}}+1\right)}{N}\,,$
and therefore with probability of at least $1-\eta$ we have that
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\epsilon_{\eta,\delta}\right)\geq\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\tilde{\epsilon}_{T}\right)\,.$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\epsilon_{\eta,\delta}\right)\geq 1-\delta\,.$
Lemma A.18.
@@
Proposition A.19.
For any $\delta>0$, with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$ and $h\sim\mathcal{P}_{\mathcal{S}}$
$\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\,,$
@@
$\epsilon=\frac{\log\left(\frac{1}{\hat{p}}\right)+4\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}}\right)\right)}{N}\,.$
@@
Here we take an alternative approach to the one presented in the proof of Theorem A.14. Denote
$\tilde{\epsilon}_{T}=\frac{\log T+2\log\log\left(T+1\right)+\log{4/\delta}}{N}\,.$
Now note that from the proof of Theorem A.14,
$T\leq\frac{\log\left(2/\eta\right)}{\hat{p}}\Rightarrow\tilde{\epsilon_{T}}\leq\epsilon_{\eta,\delta}\,.$ (24)
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\epsilon_{\eta,\delta}\right)$ $=$
$=\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\epsilon_{\eta,\delta}\middle|T\leq\frac{\log\left(2/\eta\right)}{\hat{p}}\right)\mathbb{P}_{h_{t}}\left(T\leq\frac{\log 2/\eta}{\hat{p}}\right)$
$+\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\epsilon_{\eta,\delta}\middle|T>\frac{\log\left(2/\eta\right)}{\hat{p}}\right)\mathbb{P}_{h_{t}}\left(T>\frac{\log 2/\eta}{\hat{p}}\right)$
$\geq\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\epsilon_{\eta,\delta}\middle|T\leq\frac{\log\left(2/\eta\right)}{\hat{p}}\right)\mathbb{P}_{h_{t}}\left(T\leq\frac{\log 2/\eta}{\hat{p}}\right)$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{\mathcal{D}}(h)\leq\epsilon_{\frac{\delta}{2},\frac{\delta}{2}}\right)$ $\geq\left(1-\frac{\delta}{2}\right)\left(1-\frac{\delta}{2}\right)$
@@
Applying Lemma A.18 multiple times we get
$\epsilon_{\frac{\delta}{2},\frac{\delta}{2}}$ $=\frac{\log\left(\frac{1}{\hat{p}}\right)+\log\left(\frac{8}{\delta}\right)+\log\left(\log\left(\frac{4}{\delta}\right)\right)+2\log\log\left(\frac{\log\left(\frac{4}{\delta}\right)}{\hat{p}}+1\right)}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}}\right)+2\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{\log\left(\frac{4}{\delta}\right)+1}{\hat{p}}\right)\right)}{N}$
$=\frac{\log\left(\frac{1}{\hat{p}}\right)+2\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}}\right)+\log\left(\log\left(\frac{4e}{\delta}\right)\right)\right)}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}}\right)+2\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}}\right)\right)+\log\left(\log\left(\log\left(\frac{4e}{\delta}\right)\right)+1\right)}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}}\right)+4\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}}\right)\right)}{N}$
Appendix B Proofs for Generalization of Quantized Neural Networks (Section 5)
@@
  • In Appendix B.1 we derive the upper bound for vanilla fully connected networks as stated in Theorem 5.3.
  • In Appendix B.2 we derive the upper bound for Neuron-Scaled fully-connected networks as stated in Theorem 5.3.
  • In Appendix B.3 we derive the upper bound for convolutional neural networks.
  • In Appendix B.4 we derive the upper bound for Channel-Scaled convolutional neural networks.
B.1 Vanilla Fully-Connected Neural Networks
@@
Theorem B.3.
For any activation function such that $\sigma(0)=0$, depth $L$, $Q$-quantized teacher with widths $D^{*}$, student with widths $D>D^{*}$ and prior $\mathcal{P}$ uniform over $Q$-quantized parametrizations, we have for Vanilla Fully Connected Networks that:
@@
where we defined $d_{0}^{\star}\triangleq d_{0}$. And by Theorem 3.2, $N=(\tilde{C}+3\log 2/\delta)/\epsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\epsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
are the weights and biases of $h_{\bm{\theta}^{\star}}$ respectively. Define
=$\mathcal{E}=\left\{h_{\bm{\theta}}\in\mathcal{H}_{D}^{\text{FC}}\middle|\forall l=1,\dots,L\;\mathbf{W}_{11}^{\left(l\right)}=\mathbf{W}^{\star\left(l\right)},\;\mathbf{W}_{12}^{\left(l\right)}=\mathbf{0}_{d_{l}^{\star}\times\left(d_{l-1}-d_{l-1}^{\star}\right)},\;\mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)}\right\}\,.$
We claim that any $h_{\bm{\theta}}\in\mathcal{E}$ is TE. To prove this, we show by induction over the layer $l$, that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$,
@@
Let $h_{\bm{\theta}}\in\mathcal{E}$. Begin from the base case, $l=1$. Since $d_{0}=d_{0}^{\star}$,
@@
so using the notation from Def. 5.1 we find that
@@
Since $h_{\bm{\theta}}\in\mathcal{E}$, we have that $\mathbf{W}_{11}^{\left(1\right)}=\mathbf{W}^{\star\left(1\right)}$ and $\mathbf{b}_{1}^{\left(1\right)}=\mathbf{b}^{\star\left(1\right)}$, so from (31) the coordinate projection is
@@
Next, assume that $\pi_{l-1}\left(f_{D}^{\left(l-1\right)}\left(\mathbf{x}\right)\right)=f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)$ for some $l\leq L-1$. Since $h_{\bm{\theta}}\in\mathcal{E}$, for any $l\in\left[L\right]$ we have that $\mathbf{W}_{12}^{\left(l\right)}=\mathbf{0}$. Therefore, following a similar argument we get
@@
For the last layer, $l=L$, the proof is identical, except for the application of the activation function $\sigma$ at the end, so an analogue to (33) is enough. Since we assume that $h_{\bm{\theta}}$ and $h_{\bm{\theta}^{\star}}$ have the same output dimension $d_{L}$, this proves that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$
$h_{\bm{\theta}}\left(\mathbf{x}\right)=\pi_{L}\left(h_{\bm{\theta}}\left(\mathbf{x}\right)\right)=h_{\bm{\theta}^{\star}}\left(\mathbf{x}\right)\,.$
That is, $\mathcal{E}\subseteq\left\{h_{\bm{\theta}}\middle|h_{\bm{\theta}}\equiv h_{\bm{\theta}^{\star}}\right\}$ and therefore
$\mathbb{P}\left(\mathcal{E}\right)\leq\mathbb{P}\left(h_{\bm{\theta}}\equiv h_{\bm{\theta}^{\star}}\right)\,.$
Finally, to calculate the probability to sample $h_{\bm{\theta}}$ in $\mathcal{E}$ we count the number of constrained parameters - parameters which are either determined by $h_{\bm{\theta}^{\star}}$ or are 00 in $\mathcal{E}$. Looking at the dimensions of $\mathbf{W}_{11}^{\left(l\right)}$, $\mathbf{W}_{12}^{\left(l\right)}$ and $\mathbf{b}_{1}^{\left(l\right)}$, we deduce that there are exactly
@@
B.2 Neuron-Scaled Fully-Connected Neural Networks
@@
Definition B.4 (Scaled-neuron FC restated).
For a depth $L$, widths $D=\left(d_{1},\dots,d_{L}\right)$, and activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, a scaled neuron fully connected neural network is a mapping $\bm{\theta}\mapsto h^{FC}_{\bm{\theta}}$ from parameters
$\bm{\theta}=\left(\left\{\mathbf{W}^{\left(l\right)}\right\}_{l=1}^{L},\left\{\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L},\left\{\bm{\gamma}^{\left(l\right)}\right\}_{l=1}^{L}\right)\,,$
@@
$\mathbf{W}^{\left(l\right)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{d_{l}},\bm{\gamma}^{\left(l\right)}\in\mathbb{R}^{d_{l}}\,,$
@@
$\forall l\in\left[L-1\right]\;f^{(l)}\left(\mathbf{x}\right)$ $=\sigma\left(\bm{\gamma}^{\left(l\right)}\odot\left(\mathbf{W}^{\left(l\right)}f^{(l-1)}\left(\mathbf{x}\right)\right)+\mathbf{b}^{\left(l\right)}\right)$
$h_{\bm{\theta}}^{FC}\left(\mathbf{x}\right)$ $=\operatorname{sign}\left(\mathbf{W}^{\left(L\right)}f^{(L-1)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(L\right)}\right).$
@@
Theorem B.6.
For any activation function such that $\sigma(0)=0$, depth $L$, $Q$-quantized teacher with widths $D^{*}$, student with widths $D>D^{*}$ and prior $\mathcal{P}$ uniform over $Q$-quantized parametrizations, we have for Scaled Fully Connected Networks that:
@@
where we defined $d_{0}^{\star}\triangleq d_{0}$. And by Theorem 3.2, $N=(\tilde{C}+3\log 2/\delta)/\epsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\epsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
The idea is similar to the proof of Theorem B.3, only that now we can define a set $\mathcal{E}$ with even more elements, which will tighten our bound. For any $l\in\left[L\right]$, write $\bm{\gamma}$ as a blocks vector
$\bm{\gamma}=\left[\begin{array}[]{c}\bm{\gamma}_{1}^{\left(l\right)}\\ \bm{\gamma}_{2}^{\left(l\right)}\end{array}\right]\in\mathbb{R}^{d_{l}}\,,$
@@
$\bm{\gamma}_{1}^{\left(l\right)}$ $\in\mathbb{R}^{d_{l}^{\star}}\,,$
$\bm{\gamma}_{2}^{\left(l\right)}$ $\in\mathbb{R}^{d_{l}-d_{l}^{\star}}\,.$
@@
=$\mathcal{E}=\left\{h_{\bm{\theta}}\in\mathcal{H}_{D}^{\text{SFC}}\middle|\forall l=1,\dots,L\;\mathbf{W}_{11}^{\left(l\right)}=\mathbf{W}^{\star\left(l\right)},\;\mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)},\;\mathbf{b}_{2}^{\left(l\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}},\;\bm{\gamma}_{1}^{\left(l\right)}=\bm{\gamma}^{\star\left(l\right)},\bm{\gamma}_{2}^{\left(l\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}}\right\}\,.$
As in the proof of Theorem B.3, we claim that any $h_{\bm{\theta}}\in\mathcal{E}$ is TE. This time, we specifically show by induction over the layer $l$, that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$,
@@
Let $h_{\bm{\theta}}\in\mathcal{E}$. Begin from the base case, $l=1$. Since $d_{0}=d_{0}^{\star}$,

### v2

@@ B.7 Proofs Using Nonuniform Learnability
[[[[of Theorem B.18]]]] From Lemma B.20 we have that
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\tilde{\varepsilon}_{T}\right)\geq 1-\delta\,.$ (24)
We can now use Lemma B.21 to obtain with probability of at least $1-\eta$ over the sampling of $h_{t}$ that $T\leq\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}$ whence
$\tilde{\varepsilon}_{T}$ $=\frac{\log T+2\log\log\left(T+1\right)+\log{4/\delta}}{N}$
$\leq\frac{\log\left(\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\right)+2\log\log\left(\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}+1\right)+\log{4/\delta}}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+\log\left(\frac{4}{\delta}\right)+\log\left(\log\left(\frac{2}{\eta}\right)\right)+2\log\log\left(\frac{\log\left(\frac{2}{\eta}\right)}{\hat{p}_{\mathcal{S}}}+1\right)}{N}\,,$
so with probability of at least $1-\eta$ we have that
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon_{\eta,\delta}(\mathcal{S})\right)\geq\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\tilde{\varepsilon}_{T}\right)\,.$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon_{\eta,\delta}(\mathcal{S})\right)\geq 1-\delta\,.$
Lemma B.22.
@@
Proposition B.23.
For any $\delta>0$, with probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and $h\sim\mathcal{P}_{\mathcal{S}}$
$\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\,,$
@@
$\varepsilon=\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+4\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)\right)}{N}\,.$
@@
Here we take an alternative approach to the one presented in the proof of Theorem B.18. Denote
$\tilde{\varepsilon}_{T}=\frac{\log T+2\log\log\left(T+1\right)+\log{4/\delta}}{N}\,.$
Now note that from the proof of Theorem B.18,
$T\leq\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\Rightarrow\tilde{\varepsilon_{T}}\leq\varepsilon_{\eta,\delta}(\mathcal{S})\,.$ (25)
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon_{\eta,\delta}(\mathcal{S})\right)$ $=$
$=\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon_{\eta,\delta}(\mathcal{S})\,\middle|\,T\leq\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\right)\mathbb{P}_{h_{t}}\left(T\leq\frac{\log 2/\eta}{\hat{p}_{\mathcal{S}}}\right)$
$+\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon_{\eta,\delta}(\mathcal{S})\,\middle|\,T>\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\right)\mathbb{P}_{h_{t}}\left(T>\frac{\log 2/\eta}{\hat{p}_{\mathcal{S}}}\right)$
$\geq\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon_{\eta,\delta}(\mathcal{S})\,\middle|\,T\leq\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\right)\mathbb{P}_{h_{t}}\left(T\leq\frac{\log 2/\eta}{\hat{p}_{\mathcal{S}}}\right)$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon_{\frac{\delta}{2},\frac{\delta}{2}}(\mathcal{S})\right)$ $\geq\left(1-\frac{\delta}{2}\right)\left(1-\frac{\delta}{2}\right)$
@@
Applying Lemma B.22 multiple times we get
$\varepsilon_{\frac{\delta}{2},\frac{\delta}{2}}(\mathcal{S})$ $=\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+\log\left(\frac{8}{\delta}\right)+\log\left(\log\left(\frac{4}{\delta}\right)\right)+2\log\log\left(\frac{\log\left(\frac{4}{\delta}\right)}{\hat{p}_{\mathcal{S}}}+1\right)}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+2\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{\log\left(\frac{4}{\delta}\right)+1}{\hat{p}_{\mathcal{S}}}\right)\right)}{N}$
$=\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+2\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+\log\left(\log\left(\frac{4e}{\delta}\right)\right)\right)}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+2\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)\right)+\log\left(\log\left(\log\left(\frac{4e}{\delta}\right)\right)+1\right)}{N}$
$\leq\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+4\log\left(\frac{8}{\delta}\right)+2\log\left(\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)\right)}{N}$
Appendix C Proofs for Generalization of Quantized Neural Networks (Section 4)
@@
  • In Appendix C.1 we derive the upper bound for vanilla fully connected networks as stated in Theorem 4.3.
  • In Appendix C.2 we derive the upper bound for Neuron-Scaled fully-connected networks as stated in Theorem 4.3.
  • In Appendix C.3 we derive the upper bound for convolutional neural networks.
  • In Appendix C.4 we derive the upper bound for Channel-Scaled convolutional neural networks.
C.1 Vanilla Fully-Connected Neural Networks
@@
Theorem C.3.
For any activation function such that $\sigma(0)=0$, depth $L$, $Q$-quantized teacher with widths $D^{*}$, student with widths $D>D^{*}$ and prior $\mathcal{P}$ uniform over $Q$-quantized parameterizations, we have for Vanilla Fully Connected Networks that:
@@
where we defined $d_{0}^{\star}\triangleq d_{0}$. And by Lemma 3.2, $N=(\tilde{C}+3\log 2/\delta)/\varepsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\varepsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
are the weights and biases of $h_{\boldsymbol{\theta}^{\star}}$ respectively. Define
=$\mathcal{E}=\left\{h_{\boldsymbol{\theta}}\in\mathcal{H}_{D}^{\text{FC}}\middle|\forall l=1,\dots,L\;\mathbf{W}_{11}^{\left(l\right)}=\mathbf{W}^{\star\left(l\right)},\;\mathbf{W}_{12}^{\left(l\right)}=\mathbf{0}_{d_{l}^{\star}\times\left(d_{l-1}-d_{l-1}^{\star}\right)},\;\mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)}\right\}\,.$
We claim that any $h_{\boldsymbol{\theta}}\in\mathcal{E}$ is TE. To prove this, we show by induction over the layer $l$, that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$,
@@
Let $h_{\boldsymbol{\theta}}\in\mathcal{E}$. Begin from the base case, $l=1$. Since $d_{0}=d_{0}^{\star}$,
@@
so using the notation from Def. 4.1 we find that
@@
Since $h_{\boldsymbol{\theta}}\in\mathcal{E}$, we have that $\mathbf{W}_{11}^{\left(1\right)}=\mathbf{W}^{\star\left(1\right)}$ and $\mathbf{b}_{1}^{\left(1\right)}=\mathbf{b}^{\star\left(1\right)}$, so from (32) the coordinate projection is
@@
Next, assume that $\pi_{l-1}\left(f_{D}^{\left(l-1\right)}\left(\mathbf{x}\right)\right)=f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)$ for some $l\leq L-1$. Since $h_{\boldsymbol{\theta}}\in\mathcal{E}$, for any $l\in\left[L\right]$ we have that $\mathbf{W}_{12}^{\left(l\right)}=\mathbf{0}$. Therefore, following a similar argument we get
@@
For the last layer, $l=L$, the proof is identical, except for the application of the activation function $\sigma$ at the end, so an analogue to (34) is enough. Since we assume that $h_{\boldsymbol{\theta}}$ and $h_{\boldsymbol{\theta}^{\star}}$ have the same output dimension $d_{L}$, this proves that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$
$h_{\boldsymbol{\theta}}\left(\mathbf{x}\right)=\pi_{L}\left(h_{\boldsymbol{\theta}}\left(\mathbf{x}\right)\right)=h_{\boldsymbol{\theta}^{\star}}\left(\mathbf{x}\right)\,.$
That is, $\mathcal{E}\subseteq\left\{h_{\boldsymbol{\theta}}\middle|h_{\boldsymbol{\theta}}\equiv h_{\boldsymbol{\theta}^{\star}}\right\}$ and therefore
$\mathbb{P}\left(\mathcal{E}\right)\leq\mathbb{P}\left(h_{\boldsymbol{\theta}}\equiv h_{\boldsymbol{\theta}^{\star}}\right)\,.$
Finally, to calculate the probability to sample $h_{\boldsymbol{\theta}}$ in $\mathcal{E}$ we count the number of constrained parameters - parameters which are either determined by $h_{\boldsymbol{\theta}^{\star}}$ or are 00 in $\mathcal{E}$. Looking at the dimensions of $\mathbf{W}_{11}^{\left(l\right)}$, $\mathbf{W}_{12}^{\left(l\right)}$ and $\mathbf{b}_{1}^{\left(l\right)}$, we deduce that there are exactly
@@
C.2 Neuron-Scaled Fully-Connected Neural Networks
@@
Definition C.4 (Scaled-neuron FC restated).
For a depth $L$, widths $D=\left(d_{1},\dots,d_{L}\right)$, and activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, a scaled neuron fully connected neural network is a mapping $\boldsymbol{\theta}\mapsto h^{FC}_{\boldsymbol{\theta}}$ from parameters
$\boldsymbol{\theta}=\left(\left\{\mathbf{W}^{\left(l\right)}\right\}_{l=1}^{L},\left\{\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L},\left\{\boldsymbol{\gamma}^{\left(l\right)}\right\}_{l=1}^{L}\right)\,,$
@@
$\mathbf{W}^{\left(l\right)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{d_{l}},\boldsymbol{\gamma}^{\left(l\right)}\in\mathbb{R}^{d_{l}}\,,$
@@
$\forall l\in\left[L-1\right]\;f^{(l)}\left(\mathbf{x}\right)$ $=\sigma\left(\boldsymbol{\gamma}^{\left(l\right)}\odot\left(\mathbf{W}^{\left(l\right)}f^{(l-1)}\left(\mathbf{x}\right)\right)+\mathbf{b}^{\left(l\right)}\right)$
$h_{\boldsymbol{\theta}}^{FC}\left(\mathbf{x}\right)$ $=\operatorname{sign}\left(\mathbf{W}^{\left(L\right)}f^{(L-1)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(L\right)}\right).$
@@
Theorem C.6.
For any activation function such that $\sigma(0)=0$, depth $L$, $Q$-quantized teacher with widths $D^{*}$, student with widths $D>D^{*}$ and prior $\mathcal{P}$ uniform over $Q$-quantized parameterizations, we have for Scaled Fully Connected Networks that:
@@
where we defined $d_{0}^{\star}\triangleq d_{0}$. And by Lemma 3.2, $N=(\tilde{C}+3\log 2/\delta)/\varepsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\varepsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
The idea is similar to the proof of Theorem C.3, only that now we can define a set $\mathcal{E}$ with even more elements, which will tighten our bound. For any $l\in\left[L\right]$, write $\boldsymbol{\gamma}$ as a blocks vector
$\boldsymbol{\gamma}=\left[\begin{array}[]{c}\boldsymbol{\gamma}_{1}^{\left(l\right)}\\ \boldsymbol{\gamma}_{2}^{\left(l\right)}\end{array}\right]\in\mathbb{R}^{d_{l}}\,,$
@@
$\boldsymbol{\gamma}_{1}^{\left(l\right)}$ $\in\mathbb{R}^{d_{l}^{\star}}\,,$
$\boldsymbol{\gamma}_{2}^{\left(l\right)}$ $\in\mathbb{R}^{d_{l}-d_{l}^{\star}}\,.$
@@
=$\mathcal{E}=\left\{h_{\boldsymbol{\theta}}\in\mathcal{H}_{D}^{\text{SFC}}\middle|\forall l=1,\dots,L\;\mathbf{W}_{11}^{\left(l\right)}=\mathbf{W}^{\star\left(l\right)},\;\mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)},\;\mathbf{b}_{2}^{\left(l\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}},\;\boldsymbol{\gamma}_{1}^{\left(l\right)}=\boldsymbol{\gamma}^{\star\left(l\right)},\boldsymbol{\gamma}_{2}^{\left(l\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}}\right\}\,.$
As in the proof of Theorem C.3, we claim that any $h_{\boldsymbol{\theta}}\in\mathcal{E}$ is TE. This time, we specifically show by induction over the layer $l$, that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$,
@@
Let $h_{\boldsymbol{\theta}}\in\mathcal{E}$. Begin from the base case, $l=1$. Since $d_{0}=d_{0}^{\star}$,